# Context Cascade Compression: Exploring the Upper Limits of Text Compression

### v1

@@ 1 Introduction
We revisit the entire pipeline of DeepSeek-OCR and found that when using the visual modality as a compression medium for text, As shown in figure 2. the LLM process shifts from $Text\rightarrow TextTokens\rightarrow LLM$ to $Text\rightarrow Image\rightarrow VisualLatentTokens\rightarrow LLM$. In this new pipeline, we hypothesize that information compression may not stem primarily from $Text\rightarrow Image$ conversion, but rather from the fact that latent tokens are a more efficient information representation compared to discrete text tokens. Based on this hypothesis, we propose a novel context compression paradigm: Context Cascade Compression (C3). In C3, we directly cascade two LLMs of different sizes. The first, smaller LLM compresses discrete text tokens into textual latent tokens, functioning similarly to a visual encoder in a VLM. The second, larger LLM then performs the downstream task. (For this paper, we focus on information preservation, tasking the model to reconstruct the original text from the compressed latent tokens, a process equivalent to the OCR task of recognizing text from visual latent tokens). Compared to DeepSeek-OCR, C3 employs a more streamlined pipeline for pure text compression, as it disregards other factors in optical character compression, such as layout, color, and information loss from the visual encoder. Consequently, C3 achieves a compression rate far exceeding that of optical methods while maintaining the same decoding accuracy. It also provides a potential theoretical upper bound for the ratio of text tokens to visual tokens in future work on optical character compression and OCR.
@@ 2 Related Work
Retrieval-Augmented Generation (RAG): RAG  (Lewis et al., 2021) adopts a different strategy by storing long contexts in an external knowledge base (e.g., a vector database) and retrieving only the most relevant snippets to inject into the model’s input at inference time. This approach excels in information-retrieval tasks but is inherently a "lossy" compression method, as it relies on the accuracy of the retrieval step and may lose global context or inter-snippet relationships.
@@
The most recent representative work in this area is DeepSeek-OCR (Wei et al., 2025), Glyph (Cheng et al., 2025). This model employs the OCR task as a bridge between vision and language, compressing long text by rendering it into images. It achieved a token compression rate of up to 10x while maintaining high decoding accuracy, proving the feasibility of the optical compression pathway. Our work, C3, offers new insights and improvements upon this paradigm.
Figure 3: An overview of the C3. C3 utilizes a cascaded two-LLM design. A smaller encoder LLM compresses a variable-length Input long Context into a fixed-length Latent Context guided by learnable Context Query tokens. Subsequently, a larger decoder LLM uses this compact Latent Context and a prompt to perform the downstream task, such as reconstructing the original text.
@@ 3.2 Context Compression Encoder LLM
The input sequence fed to the encoder is a concatenation of the original long context (as text tokens) context query embeddings. Crucially, the model processes this hybrid sequence uniformly. The context query embeddings are treated identically to standard text tokens within the model’s self-attention mechanism. No architectural modifications, such as introducing cross-attention layers, are required. The entire forward pass relies solely on the model’s native causal attention mechanism. This design choice ensures simplicity and leverages the full expressive power of the pre-trained Transformer architecture.Upon completion of the forward pass, the final layer’s output hidden states corresponding to the positions of the context query tokens are extracted. This resulting tensor, of shape $N\times D$, serves as an efficient and dense representation of the original text. This output constitutes the latent context, which is then passed to the downstream decoder.
@@ 3.3 Decoder LLM
For the scope of this paper, our investigation is focused exclusively on the fundamental task of text reconstruction. This task serves as a direct and rigorous benchmark for evaluating the information fidelity of our C3 compression architecture. By tasking the model with perfectly recreating the original input text from its compressed representation, we can quantitatively measure the degree of information preserved throughout the compression-decompression cycle. This setup provides a clear, objective measure of the compression’s "lossiness" and establishes a baseline for the model’s capabilities.
@@ 4.3.1 Quantitative
The superior performance of C3 can be attributed to its fundamental architectural design. The Deepseek-OCR analysis hypothesizes that its performance decline is due to factors like "complex layout" and "image blurring at lower resolutions"—inherent limitations of the optical pathway. Our C3 paradigm, by operating directly in the textual domain, is entirely immune to these visual-domain artifacts. It avoids the information loss associated with rendering text to pixels and then encoding those pixels. Instead, it leverages a pre-trained LLM’s powerful semantic understanding to distill textual information directly into an efficient latent representation.
@@ 4.3.2 Qualitative
More notably, the bottom two examples test the model’s robustness against unconventional inputs: an English paragraph interspersed with non-semantic random characters, and a structurally disordered Chinese paragraph. In both challenging scenarios, C3 successfully reconstructs the input verbatim, including the random characters and the disordered sentence structure. It effectively learns to "zip" and "unzip" the textual data, regardless of its semantic coherence.
In addition to these successful cases, it is instructive to examine the model’s failure modes under extreme pressure. When errors do occur, they tend to be concentrated in the latter of the text. In contrast to optical compression methods, which exhibit a uniform degradation of information across the entire text as the number of latent tokens decreases, our approach demonstrates a sequential information loss. Fidelity is often perfectly preserved for the initial parts of the text, while errors gradually manifest towards the end. This behavior is more analogous to the human process of forgetting, where information loss occurs progressively over time rather than as a diffuse, simultaneous blur. As show in Figure 5 This suggests that while the fixed-length latent context captures the vast majority of information, the immense compression pressure may lead to a gradual decay of fidelity for the final portions of the source text.
Figure 4: Qualitative results of text reconstruction using C3 at an extreme compression level (32 latent tokens). Each panel displays the original long text on the left and the model’s reconstructed output on the right. The examples showcase the model’s high-fidelity performance across diverse scenarios, including: (a) standard English prose, (b) classical Chinese, (c) English text containing non-semantic random characters, and (d) structurally disordered Chinese text. The near-perfect reconstruction in all cases highlights C3’s capability for near-lossless compression. Furthermore, we present an analysis of failure cases that occur under extreme compression ratios. (e) A key observation is that in these "bad cases," reconstruction errors tend to be concentrated in the latter half of the original text.
@@ 5.2 Future Work
Enabling Ultra-long Context for LLMs: The most immediate application of C3 is to serve as a powerful "front-end" compressor for existing LLMs. By compressing context that exceeds a model’s native window size (e.g., millions of tokens) into a manageable number of latent tokens, C3 can unlock new capabilities in processing entire books, extensive legal documents, or large codebases for tasks like question-answering, summarization, and analysis.This will significantly reduce costs.
Multimodal Cascaded Architecture: Reducing Costs in Long-Context Scenarios by Integrating a Lightweight VLM with an LLM for Processing Visually-Rich Documents and Multi-Image Inputs.
A Foundational Component for Next-Generation Generative Models: C3 can function as an autoencoder architecture for encoding and decoding variable-length text. The resulting latent tokens are highly applicable to next-generation generative models, such as Diffusion Language Models (DLMs) and Latent Auto-regressive Models.

### v2

@@ 1 Introduction
We revisit the entire pipeline of DeepSeek-OCR and find that when using the visual modality as a compression medium for text, as shown in figure 2, the LLM process shifts from $Text\rightarrow TextTokens\rightarrow LLM$ to $Text\rightarrow Image\rightarrow VisualLatentTokens\rightarrow LLM$. In this new pipeline, we hypothesize that information compression may not stem primarily from $Text\rightarrow Image$ conversion, but rather from the fact that latent tokens are a more efficient information representation compared to discrete text tokens. Based on this hypothesis, we propose a novel context compression paradigm: Context Cascade Compression (C3). In C3, we directly cascade two LLMs of different sizes. The first, smaller LLM compresses discrete text tokens into textual latent tokens, functioning similarly to a visual encoder in a VLM. The second, larger LLM then performs the downstream task. (For this paper, we focus on information preservation, tasking the model to reconstruct the original text from the compressed latent tokens, a process equivalent to the OCR task of recognizing text from visual latent tokens). Compared to DeepSeek-OCR, C3 employs a more streamlined pipeline for pure text compression, as it disregards other factors in optical character compression, such as layout, color, and information loss from the visual encoder. Consequently, C3 achieves a compression rate far exceeding that of optical methods while maintaining the same decoding accuracy. It also provides a potential theoretical upper bound for the ratio of text tokens to visual tokens in future work on optical character compression and OCR.
@@ 2 Related Work
Retrieval-Augmented Generation: RAG (Lewis et al., 2021) adopts a different strategy by storing long contexts in an external knowledge base (e.g., a vector database) and retrieving only the most relevant snippets to inject into the model’s input at inference time. This approach excels in information-retrieval tasks but is inherently a “loss” compression method, as it relies on the accuracy of the retrieval step and may lose global context or inter-snippet relationships.
@@
The most recent representative work in this area are DeepSeek-OCR (Wei et al., 2025), Glyph (Cheng et al., 2025). This model employs the OCR task as a bridge between vision and language, compressing long text by rendering it into images. It achieves a token compression rate of up to 10x while maintaining high decoding accuracy, proving the feasibility of the optical compression pathway. Our work, C3, offers new insights and improvements upon this paradigm.
Figure 3: An overview of the C3, which utilizes a cascaded two-LLM design. A smaller encoder LLM compresses a variable-length Input Long Context into a fixed-length Latent Context guided by learnable Context Query tokens. Subsequently, a larger decoder LLM uses this compact Latent Context and a prompt to perform the downstream task, such as reconstructing the original text.
@@ 3.2 Context Compression Encoder LLM
The input sequence fed to the encoder is a concatenation of the original long context (as text tokens) and context query embeddings. Crucially, the model processes this hybrid sequence uniformly. The context query embeddings are treated identically to standard text tokens within the model’s self-attention mechanism. No architectural modifications, such as introducing cross-attention layers, are required. The entire forward pass relies solely on the model’s native causal attention mechanism. This design choice ensures simplicity and leverages the full expressive power of the pre-trained Transformer architecture.Upon completion of the forward pass, the final layer’s output hidden states corresponding to the positions of the context query tokens are extracted. This resulting tensor, of shape $N\times D$, serves as an efficient and dense representation of the original text. This output constitutes the latent context, which is then passed to the downstream decoder.
@@ 3.3 Decoder LLM
For the scope of this paper, our investigation is focused exclusively on the fundamental task of text reconstruction. This task serves as a direct and rigorous benchmark for evaluating the information fidelity of our C3 compression architecture. By tasking the model with perfectly recreating the original input text from its compressed representation, we can quantitatively measure the degree of information preserved throughout the compression-decompression cycle. This setup provides a clear, objective measure of the compression’s “lossiness” and establishes a baseline for the model’s capabilities.
@@ 4.3.1 Quantitative
The superior performance of C3 can be attributed to its fundamental architectural design. The Deepseek-OCR analysis hypothesizes that its performance decline is due to factors like “complex layout” and “image blurring at lower resolutions”—inherent limitations of the optical pathway. Our C3 paradigm, by operating directly in the textual domain, is entirely immune to these visual-domain artifacts. It avoids the information loss associated with rendering text to pixels and then encoding those pixels. Instead, it leverages a pre-trained LLM’s powerful semantic understanding to distill textual information directly into an efficient latent representation.
@@ 4.3.2 Qualitative
More notably, the bottom two examples test the model’s robustness against unconventional inputs: an English paragraph interspersed with non-semantic random characters, and a structurally disordered Chinese paragraph. In both challenging scenarios, C3 successfully reconstructs the input verbatim, including the random characters and the disordered sentence structure. It effectively learns to “zip” and “unzip” the textual data, regardless of its semantic coherence.
In addition to these successful cases, it is instructive to examine the model’s failure modes under extreme pressure. When errors do occur, they tend to be concentrated in the latter of the text. In contrast to optical compression methods, which exhibit a uniform degradation of information across the entire text as the number of latent tokens decreases, our approach demonstrates a sequential information loss. Fidelity is often perfectly preserved for the initial parts of the text, while errors gradually manifest towards the end. This behavior is more analogous to the human process of forgetting, where information loss occurs progressively over time rather than as a diffuse, simultaneous blur. As show in Figure 5, this suggests that while the fixed-length latent context captures the vast majority of information, the immense compression pressure may lead to a gradual decay of fidelity for the final portions of the source text.
Figure 4: Qualitative results of text reconstruction using C3 at an extreme compression level (32 latent tokens). Each panel displays the original long text on the left and the model’s reconstructed output on the right. The examples showcase the model’s high-fidelity performance across diverse scenarios, including: (a) standard English prose, (b) classical Chinese, (c) English text containing non-semantic random characters, and (d) structurally disordered Chinese text. The near-perfect reconstruction in all cases highlights C3’s capability for near-lossless compression. Furthermore, we present an analysis of failure cases that occur under extreme compression ratios. (e) A key observation is that in these “bad cases”, reconstruction errors tend to be concentrated in the latter half of the original text.
@@ 5.2 Future Work
Enabling Ultra-long Context for LLMs: The most immediate application of C3 is to serve as a powerful “front-end” compressor for existing LLMs. By compressing context that exceeds a model’s native window size (e.g., millions of tokens) into a manageable number of latent tokens, C3 can unlock new capabilities in processing entire books, extensive legal documents, or large codebases for tasks like question-answering, summarization, and analysis.This will significantly reduce costs.
Multimodal Cascaded Architecture: Reducing costs in long-context scenarios by integrating a lightweight VLM with a LLM for processing visually-rich Documents, multi-image and video inputs.
A Foundational Component for Next-Generation Generative Models: C3 can function as an autoencoder architecture for encoding and decoding variable-length text. The resulting latent tokens are highly applicable to next-generation generative models, such as Diffusion Language Models and Latent Auto-Regressive Models.